OpenStudy (anonymous):

f(x) = -3 e^x + 2 + e^-5 find f'(0)

zepdrix (zepdrix):

\[\Large\bf\sf f(x)\quad=\quad-3e^{x}+2+e^{-5}\]Have you tried finding f'(x) yet?

zepdrix (zepdrix):

Did I type that correctly? Or is the +2 supposed to be in the exponent also?

OpenStudy (anonymous):

-3 e^(x+2) is f'(x)

OpenStudy (anonymous):

yeah +2 supposed to be in the exponent too

OpenStudy (anonymous):

\[f(x)=-3e ^{x+2}+e ^{-5}=-3e ^{2}e ^{x}+e ^{-5}\] \[f'(x)=-3e ^{2}e ^{x},put ~x=0~and~get~the~result.\]

OpenStudy (anonymous):

thanks, i got it :D

OpenStudy (anonymous):

Whenever you come across function notation that tells you to find f(4) (or something), always plug that into X. F(x)=1x^2+12 F(4)=1(4)^2+12 :)

OpenStudy (anonymous):

got chu

zepdrix (zepdrix):

bless you.

zepdrix (zepdrix):

That was a sneeze right? 0_o

OpenStudy (anonymous):

LOL too funny